You are here: Foswiki>PartnerObserving Web>ItalyQueue2010>ObsInaf20100613 (22 Jun 2010, VincenzoTest)Edit Attach

INAF LBC+LUCIFER queue observing, 2010 June 12-13

Observer: V. Testa, K.Boutsia
Telescope Operator: D.Gonzales Huerta
Support Astronomer: M. Pedani
Instrument: LBC

Summary

No flats - quite cloudy.
1h50min lost due to thick clouds - dome had to be closed.
Seeing very variable throughout the night due to the clouds.
~1.7h lost due to problems with the controller.

Data Summary

Project Objects Filter ExpTime Notes
Giallongo Q1623 U-Spec, z-SLOAN 0.5h seeing >1.2", out of specs
Ferraro n5466 B-bessel, r-SLOAN 40min seeing >1.2 , out of specs
Ferraro n5053 B-bessel, r-SLOAN 26min probably out of specs
Rigliaco Serpens U-Spec, R-bessel 20min  

Details

02:32 - Bias
  • 10Bias_Bino

02:41 - Pointing for flats
  • Flat script in binocular mode SkyFlat_Br_5 (sky cloudy - flats taken mainly for training)
  • flat saturated - exposure interrupted
  • Flat script in binocular mode SkyFlat_Usz_5 (0.5sec)
  • Flat script in binocular mode SkyFlat_Br_5 (1 sec - still saturated)
  • Flat script in binocular mode SkyFlat_Br_5 (1 sec)

03:00 - Thick clouds coming in - DOME CLOSED

04:50 - OPENING DOME

04:55 - Slew to Pointing Star - ACT0260 Corrected CA -66.8 IE -43.8

05:02 - Slew to NGC 5053 and dofpia

After one good iteration red underwent coma

05:14 - Given Z8 -2000 to the red then dofpia,/RedOnly

Red mirror is 2 deg warmer than outside

05:20 - dofpia both channels
  • BLUE 492 -107 -123 -207 44 -171 36 1.33 0.26
  • RED -5 -213 -47 139 -90 172 -30 0.99 0.17

05:25 - Exposure - n5053_f2_2 (Ferraro)

05:43 - dofpia
  • BLUE -437 -217 34 37 -92 153 22 1.33 0.23
  • RED -99 62 264 172 -151 49 18 1.09 0.16

05:48 - Exposure - n5053_f2_3 (Ferraro)

06:07 - Slew to Ferraro - n5466 -dofpia
  • BLUE -139 -142 -85 -70 74 7 24 1.33 0.12
  • RED 337 -66 165 92 -12 46 -13 0.89 0.17

06:13 - Exposure - n5466_f2_4 (Ferraro)
  • Seeing around 1.3".

06:29 - dofpia
  • BLUE -100 175 -33 41 83 39 5 1.39 0.11
  • RED 274 -151 -14 6 16 -107 46 1.02 0.16

06:33 - Exposure - n5466_f2_5 (Ferraro)

06:50 - dofpia, then n5466_field
  • BLUE -761 -375 -250 -64 -108 13 74 1.57 0.34
  • RED 352 11 179 -92 -4 -80 77 1.16 0.18
  • BLUE -366 -297 -294 -165 -17 -15 9 1.41 0.22
  • RED -239 191 65 -51 -63 62 6 1.02 0.15

06:56 - Exposure - n5466_field (Ferraro)
  • Seeing around 1.2-1.3 as usual

07:13 - Slew to Q1623 and focus: dofpia
  • BLUE 683 192 509 33 -97 104 12 1.30 0.33
  • RED 1346 -155 -8 -63 251 -126 -87 0.92 0.56
  • BLUE -607 178 5 79 26 170 -8 1.14 0.28
  • RED 343 223 -3 -26 -10 -165 -26 0.93 0.20
  • BLUE 427 -130 43 11 -122 -75 0 1.27 0.20
  • RED (conv'd) 309 -187 -73 31 -60 -164 14 0.94 0.19

07:19 - Q1623_Uspec_1
  • seeing>1.2''

07:54 - dofpia
  • BLUE -506 -169 -137 -97 161 218 -8 1.13 0.28
  • RED -254 182 -27 -111 -161 -62 -4 0.86 0.17

08:01 - Blue ccd controller error. Fiber link. reinitializing failed.
  • Power off/on cycle

08:06 - camera READY - dofpia
  • BLUE -105 330 120 -59 -57 -5 -28 1.06 0.14
  • RED 302 250 -149 -157 -38 -121 -20 0.79 0.20

08:09 - Q1623_Uspec_2

08:17 - First blue image delayed due to another fiber link that happened just before the end of the exposure and that was recovered automatically. Almost 5 min delay.

08:22 - Happened again at the end of the exposure. this time the system did not recover. Power cycle again, then cut exposures in 2x150s.

08:31 - restarted the camera. dofpia again
  • BLUE -194 -205 -22 84 -62 -47 47 1.30 0.14
  • RED 752 -114 193 -74 -67 -85 66 1.02 0.32
  • BLUE -433 -201 0 -30 97 1 34 1.31 0.20
  • RED -135 268 -80 -12 -38 -76 43 1.00 0.14

08:36 - Q1623_Uspec_2 again
  • Uspec 2x150s instead of 1x300s
  • Seeing around 1.4". Useless, stop the OB and go to Rigliaco field_1

08:50 - Slew to Rigliaco Seprens field 1 and dofpia
  • RED -1137 -348 20 -303 74 111 32 0.96 0.49
  • BLUE -779 415 -149 -306 4 158 15 1.28 0.38
  • RED -6 -77 -128 -34 -101 -6 18 0.96 0.10
  • BLUE -286 -102 180 3 96 192 -15 1.27 0.20

08:56 - Exposure - Rigliaco_Serpens_UsR_1

too few downloaded pixels on the first blue and red images. Blue controller and/or fpga are definitely having problems tonight, but the red problem is quite new.... also a too many downloaded pixels from the blue

09:28 - dofpia
  • RED 327 -2 165 69 -112 -47 -41 0.84 0.17
  • BLUE 1613 240 -257 -132 -209 250 -32 1.19 0.68
  • RED -5 -60 -150 -43 -109 41 1 0.92 0.11
  • BLUE 169 -206 75 78 105 -33 -23 1.13 0.14

09:31 - Exposure - Rigliaco_Serpens_UsI_1

09:36 - Blue ccd controller error again . Power off/on again

09:41 - No way. Blue FPGA won't load. Power cycle again

Meanwhile error message lbckill/lbcstart like

09:48 - Restarted

09:52 - No way, ccd controller error again.

09:56 - Talked with Fernando, may be a fan is not working. Power off and let it cool for ~15 minutes. Then try restarting again.

10:03 - Turn OFF hanged so lbckill/lbcstart but this restarted the camera, FPGA not loaded. Power OFF then wait 15 minutes, night is practically finished.

10:18 - Power ON camera.....ok!

10:23 - Try dofpia again
  • BLUE -1615 -45 -230 -40 137 17 32 1.13 0.65
  • RED -109 -118 211 -139 -187 7 12 0.90 0.16
  • BLUE -321 45 160 23 4 58 3 1.12 0.16
  • RED -234 -71 -151 83 -99 10 11 0.90 0.14

10:27 - Serpens UsI 1

10:39 - After 8 images out of 32 CCD controller error again. Turn off and let it off until flat fields.

-- VincenzoTest - 13 Jun 2010
Topic revision: r8 - 22 Jun 2010, VincenzoTest
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback